OpenStudy (anonymous):

the fourth proportional 6, 8, 15 is..........?

Directrix (directrix):

6 is to 8 as 15 is to x where x is the fourth proportional. 6/8 = 15/x 3/4 = 15/x 3x = 4*15 x = 20

Directrix (directrix):

What this means is the following: 6 / 8 = 15 / 20